OpenStudy (anonymous):

Find the zeros of the function algebraically. f(x) = √8x + 7

OpenStudy (anonymous):

ans would be f(0)=7 right ?

OpenStudy (anonymous):

That would just give y intercept

OpenStudy (anonymous):

so what would be more ....can u explain....

OpenStudy (anonymous):

Ok. So you solve the equation in terms of X. \[\sqrt{8}x = -7\]

OpenStudy (anonymous):

Then divide both sides by the square root of 8

OpenStudy (anonymous):

\[x = -7/\sqrt{8} = -2.474873\]

OpenStudy (asnaseer):

your answer depends on whether your question is:\[1) f(x)=\sqrt{8x}+7\]or:\[2) f(x)=\sqrt{8}x+7\]i..e is 'x' inside the square root symbol or not?

OpenStudy (anonymous):

yea ....bt i have taken it 2nd one...as u written

OpenStudy (asnaseer):

@amitlpu91 - you have found f(0) rather than (what I believe the question is asking) - find the value for 'x' that makes f(x) = 0

OpenStudy (asnaseer):

@mlhuyck - can you please clarify whether your question is in form of (1) or (2) above?

OpenStudy (anonymous):

bt in this case whatever value plug in for x we can't get f(x)=0 right ...

OpenStudy (asnaseer):

if the question is of the form (s) above, then @sinusoidal has solved it correctly.

OpenStudy (asnaseer):

*(2)

OpenStudy (anonymous):

watever bt idea is jst like same ......

OpenStudy (asnaseer):

f(0) is NOT the same as f(x)=0

OpenStudy (anonymous):

yea i know...

OpenStudy (anonymous):

Evaluating the fuction with x = 0 gives the y intercept

OpenStudy (anonymous):

f(0) means that here x=o right...

OpenStudy (asnaseer):

I believe the equation is of the form (2) above, so @sinusoidal has given the correct solution.

OpenStudy (anonymous):

Yes. x = 0 is a y intercept.

OpenStudy (anonymous):

y = 0 is an x intercept.

OpenStudy (anonymous):

For the graph to intercept the x axis, y must equal zero.

OpenStudy (anonymous):

also u can write as f(x)=y .....& hence proceed

OpenStudy (anonymous):

you just wrote y = y basically

OpenStudy (anonymous):

y and f(x) are both generally the dependent variable.

OpenStudy (anonymous):

no it can be considered as f(x)=y so that we can directly plug in value of x and y =0 respectively....got it

OpenStudy (anonymous):

as given above equation...

OpenStudy (anonymous):

so bcam as f(x) = y =sqr 8x +7 now it is in simplified form ...got it